OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line parallel to y = –8x – 3 through (–1, 4)

OpenStudy (aripotta):

what do parallel lines have in common?

OpenStudy (anonymous):

y intercept

OpenStudy (anonymous):

?

OpenStudy (aripotta):

no, they share the same slope

OpenStudy (anonymous):

oh ok

OpenStudy (aripotta):

so if the equation of one line is y = –8x – 3, what's the slope of the other line?

OpenStudy (anonymous):

8

OpenStudy (aripotta):

-8. the - matters :)

OpenStudy (aripotta):

do you know what point-slope form is? because i'd recommend putting the info in point-slope form first

OpenStudy (anonymous):

y=mx+b

OpenStudy (aripotta):

y = mx + b is slope-intercept form, not point-slope. point-slope is y - y1 = m(x - x1)

OpenStudy (anonymous):

PARALLEL lines have same slope PERPENDICULAR lines, you find reciprocal( 3/4's reciprocal would be 4/3) and switch the sign (+ or - depending)

OpenStudy (anonymous):

ok

OpenStudy (aripotta):

yea, the two lines are parallel lol.

OpenStudy (aripotta):

so they have the same slope...

OpenStudy (aripotta):

anyway, in y - y1 = m(x - x1) you just have to put in your given information which is: y1 = 4, x1 = -1, and m = -8

OpenStudy (anonymous):

ooh ok i confused point slope and slope intercept form

OpenStudy (aripotta):

that's ok :) it happens to me to sometimes lol. what do you get after putting the info in?

OpenStudy (aripotta):

too*

OpenStudy (anonymous):

y-4=-8(x+1)

OpenStudy (anonymous):

becuz x-(-1) = X+1

OpenStudy (aripotta):

that's right :D do you know how to change it into slope-intercept form after that?

OpenStudy (anonymous):

not sure how to

OpenStudy (aripotta):

alright, we have y - 4 = -8(x + 1) and slope-intercept form is y = mx + b. to change it into slope-intercept, you have to distribute that -8 into the parentheses, then abb 4 to both sides

OpenStudy (aripotta):

add* :l

OpenStudy (anonymous):

so y=-8x-4

OpenStudy (aripotta):

yup :)

OpenStudy (anonymous):

thanks

OpenStudy (aripotta):

no problemo